OpenStudy (anonymous):

Predict the next number in this sequence: 1, 4, 9, 16, ... A. 25 B. 32 C. 23 the answer is c right?

OpenStudy (anonymous):

21

OpenStudy (anonymous):

Wait.

OpenStudy (anonymous):

16 + 9 = 25 So A

OpenStudy (anonymous):

4-1 = 3 9-4 = 5 16 - 9 = 7 It's going up by odds.

OpenStudy (anonymous):

1+3=4 4+5=9 9+7=16 What do you think the next one should be?

OpenStudy (anonymous):

32?

OpenStudy (anonymous):

The next number is 16+9=..

OpenStudy (anonymous):

25 of course, you're right AnwarA! :D